# MonoGSDF: Exploring Monocular Geometric Cues for Gaussian Splatting-Guided Implicit Surface Reconstruction

### v3

@@ Abstract
Accurate meshing from monocular images remains a key challenge in 3D vision. While state-of-the-art 3D Gaussian Splatting (3DGS) methods excel at synthesizing photorealistic novel views through rasterization-based rendering, their reliance on sparse, explicit primitives severely limits their ability to recover watertight and topologically consistent 3D surfaces. We introduce MonoGSDF, a novel method that couples Gaussian-based primitives with a neural Signed Distance Field (SDF) for high-quality reconstruction. During training, the SDF guides Gaussians’ spatial distribution, while at inference, Gaussians serve as priors to reconstruct surfaces, eliminating the need for memory-intensive Marching Cubes. To handle arbitrary-scale scenes, we propose a scaling strategy for robust generalization. A multi-resolution training scheme further refines details and monocular geometric cues from off-the-shelf estimators enhance reconstruction quality. Experiments on real-world datasets show MonoGSDF outperforms prior methods while maintaining efficiency.
@@ 1 Introduction
Figure 1: MonoGSDF. Ours achieves superior surface reconstruction on arbitrary-scale scenes, particularly in areas with limited view coverage. Compared to 2DGS [12] and GOF [48], ours achieves higher F1 scores ($\uparrow$) and reconstructs smooth surfaces with fine details.
@@
3D Gaussian Splatting (3DGS) [16] has emerged as a state-of-the-art method for high-quality novel view synthesis by leveraging the rasterization pipeline and representing 3D scenes using points characterized by Gaussian functions. However, 3DGS remains an explicit and discrete representation, posing challenges for accurate 3D surface reconstruction and meshing. This limitation is particularly important for applications such as geometry editing [43, 6, 36], 3D animation [33, 49, 30], and robotics [56, 15, 50]. The primary challenge lies in 3DGS being a discrete and unstructured point-based geometry representation. Given that 3D Gaussians are primarily optimized for accurate appearance rendering rather than geometric regularization, which complicates the extraction of accurate 3D surfaces through post-processing methods like Poisson surface reconstruction [14].
@@
These challenges have motivated recent investigations exploring how 3DGS can be used for high-quality surface reconstruction while maintaining its rendering speed and efficiency. Methods like 2DGS [12] address this by aligning 2D Gaussians to surfaces, although their depth map-based mesh extraction [7] struggles with thin structures and unbounded scenes due to resolution and resource limitations. Methods such as SuGaR [11] attempt to address these limitations by aligning 3D Gaussians with mesh faces, while DN-Splater [34] incorporates depth and normal estimators. Nevertheless their reliance on Poisson reconstruction [14] still fundamentally limits reconstruction and meshing quality due to the inherent geometric inaccuracies of Gaussian representations. Although recent methods [45, 24, 55] have integrated SDF fields into Gaussian training for surface reconstruction, their incomplete exploration of the Gaussian-SDF relationship results in over-smoothed surfaces and limits their application within object-level reconstruction.
To bridge the gap between explicit Gaussian representations and implicit surface reconstruction while fully leveraging their complementary advantages, we present MonoGSDF, a unified framework that establishes a robust connection between 3D Gaussians and SDF-based surface modeling for high-fidelity reconstruction. Our contributions are:
  • A Gaussian-based implicit surface reconstruction pipeline for monocular images, where the SDF guides Gaussians to better align near the surface during training, while Gaussians act as priors for fast zero-level set extraction during reconstruction.
  • A simple yet efficient SDF-to-Opacity function that bridges implicit SDF with explicit Gaussians, coupled with a Gaussian-guided normalization strategy that dynamically maps unbounded 3D coordinates to a (0,1) range with minimal distortion.
  • A multi-resolution training strategy based on wavelet transforms progressively refines details, while off-the-shelf geometric cues further enhance reconstruction quality, enabling Gaussians to better capture scene geometry.
Experiments on real-world datasets demonstrate that our method delivers high-quality reconstructions and outperforms baselines by 13% in terms of Chamfer distance while preserving the view synthesis quality.
Figure 2: Overview. MonoGSDF establishes a synergistic relationship between SDF and Gaussian representation: the SDF regulates Gaussian distribution during training, while Gaussians serve as efficient primitives for rapid surface extraction. We first query the SDF values for each Gaussians and convert them into opacity values. This enforces Gaussian to stick closer to the surface. Then we follow the standard rasterization pipeline to render color, depth and normal maps. Off-the-shelf geometry cues and multi-resolution strategy are applied for better geometry performance. And we jointly train the SDF via 3D supervision signal from rendered Gaussian depth.
@@ 2.1 View Synthesis and Gaussian Splatting
Neural Radiance Field (NeRF) [25] utilize multi-layer perceptions (MLP) as scene representation to predict geometry and view-dependent appearance. The MLP is optimized via a photometric loss through volume rendering. Subsequent methods have focused on optimizing NeRF’s training and expressiveness using grid representations [28], improving rendering speed [32, 53] and scaling to unbounded scenes [3]. However, volume rendering typically requires substantial computational resources and extensive training durations. 3D Gaussian Splatting (3DGS) [16] has emerged as an efficient approach for real-time view synthesis through differentiable Gaussian functions. Subsequent works have enhanced rendering quality via anti-aliasing techniques [47, 40, 31] and improved rendering speed through Gaussian density control [42] and radiance field priors [22, 29]. Geometry-focused approaches like DNGaussian [19] address sparse view degradation, while GeoGaussian [20] preserves non-textured regions. Instantsplat [9] accelerates sparse view training using Dust3r [38] initialization, and Scaffold-GS [23] combines implicit-explicit representations. However, these methods primarily focus on appearance quality rather than underlying geometry, limiting their application to view synthesis.
@@ 2.2 Surface Reconstruction from Gaussians
Due to the discrete and unstructured nature of 3DGS, along with supervision being limited to RGB images during training, existing methods often struggle to accurately capture scene geometry, making surface reconstruction particularly challenging. SuGaR [11] addresses this by constructing a density field from Gaussians and extracting meshes via level-set searching. However, it is computationally expensive and struggles to reconstruct large, smooth surfaces like floors and walls. Other approaches [4, 5, 51, 34, 39] incorporate depth or normal estimators as priors to supervise Gaussians. 2DGS [12] and GSurfels [8] improves geometric alignment by flattening 3D Gaussians into 2D disks or surfels, enabling better surface representation. For surface reconstruction, they rely on either multi-view depth maps for TSDF fusion or Poisson reconstruction. However, TSDF fusion is constrained by fixed resolution and high memory usage for large scenes. To handle unbounded scenes, it requires contracting coordinates, introducing unnecessary distortions. Meanwhile, Poisson reconstruction often produces noisy surfaces due to inaccuracies in Gaussian geometry distribution. Instead of refining depth and normal maps, some methods [45, 55, 24, 2] integrate Signed Distance Field (SDF). 3DGSR [24] and GSDF [45] associate Gaussians with SDF via an additional branch that volume-renders depth and normal supervised by rendered Gaussian depth and normal. However, this approach is inefficient and fails to fully exploit the relationship between Gaussians and SDF. GS-Pull [55] and GSurf [2] improve upon this by leveraging SDF gradients to better align Gaussians with surfaces, but they are either restricted to object-level reconstruction or produce overly smooth results. Gaussian Opacity Fields (GOF) [48] approximates minimum accumulated alpha values from all views via ray tracing to construct an opacity field, followed by surface extraction using Marching Tetrahedra [18]. However, Gaussian distributions remain inconsistent and unstructured relative to scene surfaces. Our method leverages SDF to guide Gaussians to align closer to the surface during training. At inference, Gaussians serve as primitives to efficiently constrain zero-level set extraction, avoiding redundant free-space searches while achieving accurate reconstruction.
@@ 3.1 Preliminaries
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) [16] employs a set of 3D points to effectively render images from given viewpoints, each characterized by a Gaussian function with 3D mean $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity value $\alpha_{i}\in\mathbb{R}$, RGB color values $\mathbf{c}_{i}\in\mathbb{R}^{3}$:
$o_{i}(\mathbf{x})=\alpha_{i}*\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})\right).$ (1)
Given a 3D position $\mathbf{x}$, $o_{i}(\mathbf{x})$ represents current opacity value contributed by the $i$-th Gaussian. To facilitate optimization, $\Sigma_{i}$ is factorized into the product of a scaling matrix $S_{i}$, represented by scale factors $\mathbf{s}_{i}\in\mathbb{R}^{3}$, and a rotation matrix $R_{i}$ encoded by a quaternion $\mathbf{r}_{i}\in\mathbb{R}^{4}$:
$\Sigma_{i}=R_{i}S_{i}S^{T}_{i}R^{T}_{i}.$ (2)
3D Gaussians are then projected onto a 2D image plane according to elliptical weighted average (EWA) [57] to render images for given views. Color $\mathbf{\bar{C}}(\mathbf{u})$, depth ${\bar{D}}(\mathbf{u})$, and normal $\mathbf{\bar{N}}(\mathbf{u})$ at pixel $\mathbf{u}$ is rendered by $N$ projected and ordered Gaussians using point-based $\alpha$-blending:
$\{\mathbf{\bar{C}},\bar{D},\mathbf{\bar{N}}\}(\mathbf{u})=\sum_{i\in N}T_{i}o_{i}\{\mathbf{c}_{i},d_{i},\mathbf{n}_{i}\},$ (3)
@@
Neural Signed Distance Field. While 3DGS excels in novel view synthesis, its discrete nature limits surface reconstruction. In contrast, SDF provide a continuous, watertight surface as the zero-level set. We use a hash-based neural representation and jointly train the SDF with 3DGS, enabling more precise and efficient surface extraction. To better encode the scene geometry, we choose One-blob $\gamma(\mathbf{x})$ [26] as positional embedding and a multi-resolution hash-based feature grid $\mathcal{V}=\{\mathcal{V}^{l}\}^{L}_{l=1}$ [28]. Feature $\mathcal{V}(h(\mathbf{x}))$ at any 3D point $\mathbf{x}$ are queried via trilinear interpolation. The geometry decoder $f$ is an MLP which maps the 3D coordinate to an SDF value $\bar{s}$:
@@ 3.2 Gaussian-Guided Signed Distance Field
Connect Gaussians with SDF. As illustrated in Figure 2, our method begins by querying SDF values with each Gaussian 3D mean $\mathbf{\mu}_{i}$ using Eq. 4, where zero values correspond to surface points. While 3DGSR [24] introduced a bell-shaped function to map SDF value to Gaussian opacity, this approach assigns an opacity value $\alpha_{i}=0.25$ to surface points ($\bar{s}=0$), which contradicts the physical reality that surface points should exhibit maximum opacity ($\alpha_{i}=1$). To address this limitation, we propose a Gaussian-shaped transformation that maintains both mathematical simplicity and computational efficiency while ensuring physically plausible opacity values $g(\bar{s})$:
$g(\bar{s}_{i})=\exp(-(\beta*\bar{s}_{i})^{2})\mapsto\alpha_{i},$ (5)
where $\beta$ is a hyperparameter. While previous 3DGS methods [16, 12, 23] optimize opacities through alpha blending without considering spatial distribution constraints, these approaches often result in limited geometric accuracy. Our method enforces a better distribution of Gaussians along the surface, leading to high geometric fidelity.
Gaussian-Guided Normalization Our rendering pipeline follows the standard Gaussian rasterization process, generating color $\mathbf{\bar{C}}$, depth $\bar{D}$ and $\mathbf{\bar{N}}$ maps from Gaussian attributes, with the depth map subsequently used for SDF training. While 3DGSR [24] and GSDF [45] have adopted similar approaches, their grid-based representations necessitate predefined scene bounding boxes for volume initialization, restricting their applicability to object-level or small-scale scene reconstruction. Although MipNeRF [3] introduced a fixed coordinate contraction method for unbounded scenes, its mapping of infinite space to [-2, 2] conflicts with our hash-based feature field’s requirement for (0, 1) input. Moreover, MipNeRF’s linear transformation within a limited region of interest introduces substantial distortion in peripheral areas. To address these limitations and to enable scalable reconstruction of real-world unbounded scenes, our method introduces a Gaussian-guided normalization strategy. We first establish a bounding box based on the Gaussians, then apply a novel sigmoid-like normalization function that maps 3D point $\mathbf{x}\in\mathbb{R}^{3}$ to a fixed $(0,1)$ range:

### v4

@@ Abstract
Accurate meshing from images remains a key challenge in 3D vision. While state-of-the-art 3D Gaussian Splatting (3DGS) methods excel at synthesizing photorealistic novel views through rasterization-based rendering, their reliance on sparse, explicit primitives severely limits their ability to recover watertight and topologically consistent 3D surfaces. We introduce MonoGSDF, a novel method that couples Gaussian-based primitives with a neural Signed Distance Field (SDF) for high-quality reconstruction with monocular RGB images as input. During training, the SDF guides Gaussians’ spatial distribution, while at inference, Gaussians serve as priors to reconstruct surfaces, eliminating the need for memory-intensive Marching Cubes. To handle arbitrary-scale scenes, we propose a scaling strategy for robust generalization. A multi-resolution training scheme further refines details and monocular geometric cues from off-the-shelf estimators enhance reconstruction quality. Experiments on real-world datasets show MonoGSDF outperforms prior methods while maintaining efficiency.
@@ 1 Introduction
Figure 1: MonoGSDF. We show the reconstructed mesh. Compared to 2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] and GOF [Yu et al.(2024c)Yu, Sattler, and Geiger], ours achieves higher F1 scores ($\uparrow$) and reconstructs smooth surfaces with fine details.
@@
3D Gaussian Splatting (3DGS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] has emerged as a state-of-the-art method for high-quality novel view synthesis by leveraging the rasterization pipeline and representing 3D scenes using points characterized by Gaussian functions. However, 3DGS remains an explicit and discrete representation, posing challenges for accurate 3D surface reconstruction and meshing. This limitation is particularly important for applications such as geometry editing [Yang et al.(2024c)Yang, Gao, Zhou, Jiao, Zhang, and Jin, Chen et al.(2024c)Chen, Chen, Zhang, Wang, Yang, Wang, Cai, Yang, Liu, and Lin, Wang et al.(2024a)Wang, Fang, Zhang, Xie, and Tian], 3D animation [Tang et al.(2023)Tang, Ren, Zhou, Liu, and Zeng, Yuan et al.(2024)Yuan, Li, Huang, De Mello, Nagano, Kautz, and Iqbal, Qian et al.(2024)Qian, Wang, Mihajlovic, Geiger, and Tang], and robotics [Zhou et al.(2024)Zhou, Lin, Shan, Wang, Sun, and Yang, Keetha et al.(2024)Keetha, Karhade, Jatavallabhula, Yang, Scherer, Ramanan, and Luiten, Yugay et al.(2023)Yugay, Li, Gevers, and Oswald]. The main challenge is that 3DGS models geometry as a discrete, unstructured point set. Because 3D Gaussians are optimized for rendering quality instead of geometric consistency, post-processing methods such as Poisson surface reconstruction [Kazhdan et al.(2006)Kazhdan, Bolitho, and Hoppe] struggle to recover accurate 3D surfaces.
@@
These challenges have motivated recent investigations exploring how 3DGS can be used for high-quality surface reconstruction while maintaining its rendering speed and efficiency. Methods like 2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] address this by aligning 2D Gaussians to surfaces, although their depth map-based mesh extraction [Curless and Levoy(1996)] struggles with thin structures and unbounded scenes due to resolution and resource limitations. Methods such as SuGaR [Guédon and Lepetit(2024)] attempt to address these limitations by aligning 3D Gaussians with mesh faces. Nevertheless their reliance on Poisson reconstruction [Kazhdan et al.(2006)Kazhdan, Bolitho, and Hoppe] still fundamentally limits reconstruction and meshing quality due to the inherent geometric inaccuracies of Gaussian representations. Although recent methods [Yu et al.(2024a)Yu, Lu, Xu, Jiang, Xiangli, and Dai, Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi, Zhang et al.(2025)Zhang, Liu, and Han] have integrated SDF fields into Gaussian training for surface reconstruction, their incomplete exploration of the Gaussian-SDF relationship results in over-smoothed surfaces and limits their application within object-level reconstruction.
To bridge the gap between explicit Gaussian representations and implicit surface reconstruction while fully leveraging their complementary advantages, we present MonoGSDF, a unified framework that establishes a robust connection between 3D Gaussians and SDF-based surface modeling for high-fidelity reconstruction with only monocular RGB images as input. Our contributions are: 1. a Gaussian-based implicit surface reconstruction pipeline for monocular images, where the SDF guides Gaussians to better align near the surface during training and Gaussians act as priors for fast mesh extraction; 2. a simple yet efficient SDF-to-Opacity function that bridges implicit SDF with explicit Gaussians, coupled with a normalization strategy that maps unbounded 3D coordinates to a (0,1) range with minimal distortion; 3. a multi-resolution strategy based on wavelet transforms progressively refines details, while off-the-shelf geometric cues further enhance reconstruction quality. Experiments on real-world datasets demonstrate that our method delivers high-quality reconstructions and outperforms baselines by 13% in terms of Chamfer distance while preserving the view synthesis quality.
@@ 2.1 View Synthesis and Gaussian Splatting
Neural Radiance Field (NeRF) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] utilize multi-layer perceptions (MLP) as scene representation to predict geometry and view-dependent appearance. The MLP is optimized via a photometric loss through volume rendering. Subsequent methods have focused on optimizing NeRF’s training and expressiveness using grid representations [Müller et al.(2022)Müller, Evans, Schied, and Keller], improving rendering speed [Sun et al.(2022)Sun, Sun, and Chen, Zhang et al.(2020)Zhang, Riegler, Snavely, and Koltun] and scaling to unbounded scenes [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]. However, volume rendering typically requires substantial computational resources and extensive training durations. 3D Gaussian Splatting (3DGS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] has emerged as an efficient approach for real-time view synthesis through differentiable Gaussian functions. Subsequent works have enhanced rendering quality via anti-aliasing techniques [Yu et al.(2024b)Yu, Chen, Huang, Sattler, and Geiger, Yan et al.(2024)Yan, Low, Chen, and Lee, Song et al.(2024)Song, Zheng, Yuan, Gao, Zhao, He, Gu, and Zhao] and improved rendering speed through Gaussian density control [Yang et al.(2024b)Yang, Zhu, Jiang, Ye, Chen, Zhang, Chen, Zhao, and Zhao] and radiance field priors [Liu et al.(2025)Liu, Wang, Hu, Shen, Ye, Zang, Cao, Li, and Liu, Niemeyer et al.(2024)Niemeyer, Manhardt, Rakotosaona, Oechsle, Duckworth, Gosula, Tateno, Bates, Kaeser, and Tombari]. Geometry-focused approaches like DNGaussian [Li et al.(2024)Li, Zhang, Bai, Zheng, Ning, Zhou, and Gu] address sparse view degradation, while GeoGaussian [Li et al.(2025)Li, Lyu, Di, Zhai, Lee, and Tombari] preserves non-textured regions. Instantsplat [Fan et al.(2024)Fan, Cong, Wen, Wang, Zhang, Ding, Xu, Ivanovic, Pavone, Pavlakos, et al.] accelerates sparse view training using Dust3r [Wang et al.(2024b)Wang, Leroy, Cabon, Chidlovskii, and Revaud] initialization, and Scaffold-GS [Lu et al.(2024)Lu, Yu, Xu, Xiangli, Wang, Lin, and Dai] combines implicit-explicit representations. However, these methods primarily focus on appearance quality rather than underlying geometry, limiting their application to view synthesis.
@@ 2.2 Surface Reconstruction from Gaussians
Due to the discrete and unstructured nature of 3DGS, along with supervision being limited to RGB images during training, existing methods often struggle to accurately capture scene geometry, making surface reconstruction particularly challenging. SuGaR [Guédon and Lepetit(2024)] addresses this by constructing a density field from Gaussians and extracting meshes via level-set searching. However, it is computationally expensive and struggles to reconstruct large, smooth surfaces like floors and walls. Other approaches [Chen et al.(2024a)Chen, Li, Ye, Wang, Xie, Zhai, Wang, Liu, Bao, and Zhang, Chen et al.(2024b)Chen, Wei, Li, Huang, Wang, and Lee, Zhang et al.(2024a)Zhang, Fang, Shrestha, Liang, Long, and Tan, Turkulainen et al.(2024)Turkulainen, Ren, Melekhov, Seiskari, Rahtu, and Kannala, Wolf et al.(2024)Wolf, Bracha, and Kimmel] incorporate depth or normal estimators as priors to supervise Gaussians. 2DGS [Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao] and GSurfels [Dai et al.(2024)Dai, Xu, Xie, Liu, Wang, and Xu] improves geometric alignment by flattening 3D Gaussians into 2D disks or surfels, enabling better surface representation. For surface reconstruction, they rely on either multi-view depth maps for TSDF fusion or Poisson reconstruction. However, TSDF fusion is constrained by fixed resolution and high memory usage for large scenes. To handle unbounded scenes, it requires contracting coordinates, introducing unnecessary distortions. Meanwhile, Poisson reconstruction often produces noisy surfaces due to inaccuracies in Gaussian geometry distribution. Instead of refining depth and normal maps, some methods [Yu et al.(2024a)Yu, Lu, Xu, Jiang, Xiangli, and Dai, Zhang et al.(2025)Zhang, Liu, and Han, Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi, Baixin et al.(2024)Baixin, Jiangbei, Jiaze, and Ying] integrate Signed Distance Field (SDF). 3DGSR [Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] and GSDF [Yu et al.(2024a)Yu, Lu, Xu, Jiang, Xiangli, and Dai] associate Gaussians with SDF via an additional branch that volume-renders depth and normal supervised by rendered Gaussian depth and normal. However, this approach is inefficient and fails to fully exploit the relationship between Gaussians and SDF. GS-Pull [Zhang et al.(2025)Zhang, Liu, and Han] and GSurf [Baixin et al.(2024)Baixin, Jiangbei, Jiaze, and Ying] improve upon this by leveraging SDF gradients to better align Gaussians with surfaces, but they are either restricted to object-level reconstruction or produce overly smooth results. Gaussian Opacity Fields (GOF) [Yu et al.(2024c)Yu, Sattler, and Geiger] approximates minimum accumulated alpha values from all views via ray tracing to construct an opacity field, followed by surface extraction using Marching Tetrahedra [Kulhanek and Sattler(2023)]. However, Gaussian distributions remain inconsistent and unstructured relative to scene surfaces. Our method leverages SDF to guide Gaussians to align closer to the surface during training. At inference, Gaussians serve as primitives to efficiently constrain zero-level set extraction, avoiding redundant free-space searches while achieving accurate reconstruction.
@@ 3.1 Preliminaries
3D Gaussian Splatting.
3D Gaussian Splatting (3DGS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] employs a set of 3D points to effectively render images from given viewpoints, each characterized by a Gaussian function with 3D mean $\mathbf{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity value $\alpha_{i}\in\mathbb{R}$, RGB color values $\mathbf{c}_{i}\in\mathbb{R}^{3}$:
$o_{i}(\mathbf{x})=\alpha_{i}\times\exp\left(-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})\right),\ \ \Sigma_{i}=R_{i}S_{i}S^{T}_{i}R^{T}_{i}.$ (1)
Given a 3D position $\mathbf{x}$, $o_{i}(\mathbf{x})$ represents current opacity value contributed by the $i$-th Gaussian. To facilitate optimization, $\Sigma_{i}$ is factorized into the product of a scaling matrix $S_{i}$, represented by scale factors $\mathbf{s}_{i}\in\mathbb{R}^{3}$, and a rotation matrix $R_{i}$ encoded by a quaternion $\mathbf{r}_{i}\in\mathbb{R}^{4}$. 3D Gaussians are then projected onto a 2D image plane according to elliptical weighted average (EWA) [Zwicker et al.(2002)Zwicker, Pfister, Van Baar, and Gross] to render images for given views. Color $\mathbf{\bar{C}}(\mathbf{u})$, depth ${\bar{D}}(\mathbf{u})$, and normal $\mathbf{\bar{N}}(\mathbf{u})$ at pixel $\mathbf{u}$ is rendered by $N$ projected and ordered Gaussians using point-based $\alpha$-blending:
$\{\mathbf{\bar{C}},\bar{D},\mathbf{\bar{N}}\}(\mathbf{u})=\sum_{i\in N}T_{i}o_{i}\{\mathbf{c}_{i},d_{i},\mathbf{n}_{i}\},\vskip-5.69046pt$ (2)
@@
Figure 2: Overview. MonoGSDF synergizes SDF and Gaussian representations by converting queried SDF values into opacity, encouraging Gaussians to align with surfaces. It uses standard rasterization to render color, depth, and normals, enhanced by geometry cues and a multi-resolution strategy. SDF is jointly trained with 3D supervision from rendered depth.
Neural Signed Distance Field.
While 3DGS excels in novel view synthesis, its discrete nature limits surface reconstruction. In contrast, SDF provide a continuous, watertight surface as the zero-level set. We use a hash-based neural representation and jointly train the SDF with 3DGS, enabling more precise and efficient surface extraction. To better encode the scene geometry, we choose One-blob $\gamma(\mathbf{x})$ [Müller et al.(2019)Müller, McWilliams, Rousselle, Gross, and Novák] as positional embedding and a multi-resolution hash-based feature grid $\mathcal{V}=\{\mathcal{V}^{l}\}^{L}_{l=1}$ [Müller et al.(2022)Müller, Evans, Schied, and Keller]. Feature $\mathcal{V}(h(\mathbf{x}))$ at any 3D point $\mathbf{x}$ are queried via trilinear interpolation. The geometry decoder $f$ is an MLP which maps the 3D coordinate to an SDF value $\bar{s}$:
@@ 3.2 Gaussian-Guided Signed Distance Field
Connect Gaussians with SDF.
As illustrated in Figure 2, our method begins by querying SDF values with each Gaussian 3D mean $\mathbf{\mu}_{i}$ using Eq. 3, where zero values correspond to surface points. While 3DGSR [Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] introduced a bell-shaped function to map SDF value to Gaussian opacity, this approach assigns an opacity value $\alpha_{i}=0.25$ to surface points ($\bar{s}=0$), which contradicts the physical reality that surface points should exhibit maximum opacity ($\alpha_{i}=1$). To address this limitation, we propose a Gaussian-shaped transformation that maintains both mathematical simplicity and computational efficiency while ensuring physically plausible opacity values $g(\bar{s})$:
$g(\bar{s}_{i})=\exp(-(\beta\times\bar{s}_{i})^{2})\mapsto\alpha_{i},$ (4)
where $\beta$ is a hyperparameter. While previous 3DGS methods [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis, Huang et al.(2024)Huang, Yu, Chen, Geiger, and Gao, Lu et al.(2024)Lu, Yu, Xu, Xiangli, Wang, Lin, and Dai] optimize opacities through alpha blending without considering spatial distribution constraints, these approaches often result in limited geometric accuracy. Our method enforces a better distribution of Gaussians along the surface, leading to high geometric fidelity.
Gaussian-Guided Normalization
Our rendering pipeline follows the standard Gaussian rasterization process, generating color $\mathbf{\bar{C}}$, depth $\bar{D}$ and $\mathbf{\bar{N}}$ maps from Gaussian attributes, with the depth map subsequently used for SDF training. While 3DGSR [Lyu et al.(2024)Lyu, Sun, Huang, Wu, Yang, Chen, Pang, and Qi] and GSDF [Yu et al.(2024a)Yu, Lu, Xu, Jiang, Xiangli, and Dai] have adopted similar approaches, their grid-based representations necessitate predefined scene bounding boxes for volume initialization, restricting their applicability to object-level or small-scale scene reconstruction. Although MipNeRF [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] introduced a fixed coordinate contraction method for unbounded scenes, its mapping of infinite space to [-2, 2] conflicts with our hash-based feature field’s requirement for (0, 1) input. Moreover, MipNeRF’s linear transformation within a limited region of interest introduces substantial distortion in peripheral areas. To address these limitations and to enable scalable reconstruction of real-world unbounded scenes, our method introduces a Gaussian-guided normalization strategy. We first establish a bounding box based on the Gaussians, then apply a novel sigmoid-like normalization function that maps 3D point $\mathbf{x}\in\mathbb{R}^{3}$ to a fixed $(0,1)$ range: